-5x+3y=51 y=10x-8 x=? y=?
Answer by TimothyLamb(4379)

---
-5x + 3y = 51
y = 10x - 8
---
put the system of linear equations into standard form:
---
-5x + 3y = 51
10x - y = 8
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3
y = 22
---
system matrix
-5 3
10 -1
---
inverse of system matrix
0.04 0.12
0.4 0.2
---
determinant of system matrix = -25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php